OpenStudy (anonymous):

Determine the solution set of the inequality -29<4x-5<11

OpenStudy (anonymous):

-29+5<4x-5+5<11+5 -24<4x<16 -24/4<4x/4<16/4 -6<x<4

OpenStudy (anonymous):

Thanks Soo Much , Could You Help Me With A Few More ?

OpenStudy (anonymous):

yes i can

OpenStudy (anonymous):

Simplify the complex fraction: 1.(2x+5y/4y) + (-2x+3y/4y) 2. Simplify the complex fraction: (1- (4/x^2))/(1+(2/x)) 3. Find the solution of the equation 3/4x-2=5/6x+2/3

OpenStudy (anonymous):

1) 2x-2x+5y+3y/4y 8y/4y=2

OpenStudy (anonymous):

Thanks Again . Would Continue And Show Me #2 and #3 ?